Two lines in a coordinate plane have no points of intersection. Which of these could be the equations of the lines?
A. 4x+2y=6 C.4x+=6
10x+5y=7 5x-10y=6
B. 4x+2y=6 D. 5x+10y=6
10x+5y=15 5x-10y=6

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53419)

.
Two lines in a coordinate plane have no points of intersection.
Which of these could be the equations of the lines?
A.  4x + 2y =  6   C. 4x +     = 6
   10x + 5y =  7      5x - 10y = 6

B.  4x + 2y =  6   D. 5x + 10y = 6
   10x + 5y = 15      5x - 10y = 6
~~~~~~~~~~~~~~~~~~~~~~

(A)  The lines have the same slope, but different y-intercepts.
     So, these lines are parallel and have no common points.


(B)  The lines have the same slope and the same y-intercepts.
     So, these lines coincide.


(C)  System (C) is written incorrectly, so I will not discuss/consider it.


(D)  The lines have different slopes, so they intersect each other.


ANSWER.  As far as I can make a conclusion considering cases (A), (B), and (D), lines (A) have no intersection points.

Solved.

Answer in the post by @lynnlo is incorrect.